# Impact of Guidance and Interaction Strategies for LLM Use on Learner Performance and Perception

### v2

@@ 2.3. Understanding Factors that Influence Learners’ Help-Seeking Decisions
To characterize how learner-LLM interactions may be shaped by important factors that influence learner help-seeking decisions more broadly, we can turn to the educational psychology literature on self-efficacy and perceived task difficulty. The choice of when to consult external tools during problem-solving isn’t exclusive to the realm of LLMs. In broader educational research, students’ decisions to access resources—be it textbooks, online forums, peers, or tutors—often come down to a balance between self-efficacy, perceived task difficulty, and the accessibility of the tool.
@@
In the context of LLMs, the aforementioned considerations take on a nuanced significance. The availability and perceived efficacy of LLMs might encourage more immediate reliance, potentially overshadowing traditional problem-solving approaches. Our paper’s exploration of different student approaches offers insights into these nuanced behaviors in the specific context of LLM interactions.
We factored such considerations into the design of our experiments and included questions that help us quantify such factors using pre- and post-measures regarding perception of LLMs (confidence in LLMs’ responses, helpfulness, willingness to interact again, and error tolerance) and self-perception (confidence in their answers, self-confidence for the given topic of the assignment). To measure long-term learning, we assessed their performance on the final exam on the same topic, comprising isomorphic questions to those in the assignment.
@@ 3. Design Considerations for Learner Interaction with LLMs
As language models continue to find their way into educational settings, understanding how students interact with LLM-based tools becomes imperative. In our exploration of learner interactions with LLMs, we identified three critical factors influencing these engagements:
  1. (1)
    Guidance Received: The directions or suggestions that students receive can shape their entire interaction experience. The form and content of this guidance can determine how students approach and utilize LLMs to their advantage (Section 3.1).
  2. (2)
    Student Approach: Students’ initial inclination, whether to consult the LLM immediately or to first wrestle with the problem on their own, can have implications for their learning process. The strategies they adopt can be influenced by the guidance they have received and can significantly vary in efficacy (Section 3.2).
  3. (3)
    LLMs’ Response: While guidance and approach set the stage, the actual dialogue with the LLM remains paramount. The quality, clarity, and relevance of the LLMs’ responses can make or break the learning experience (Section 3.3).
@@ 3.1. Guidance Strategies for LLM Engagement
We devised a set of suggestions for LLM chatbot use, as depicted in Figure 1A. Given the limited existing research on LLM guidance for students, our suggestions drew from prominent online sources (Pangu, 2023; Miquelino, [n. d.]; Mollick, 2023b). Recognizing the growing classroom trend of prescribed LLM instructions (eth, 2023; Mollick, 2023a), G1 serves as a representative model for evaluating its efficacy against other guidance strategies.
@@
Leveraging examples is an intuitive approach to tool introduction (Lee, 2004). However, the choice of examples plays a crucial role. For our classroom deployment, we took examples of interactions from previous interview studies with students [REDACTED]. These examples, validated by an instructor and a teaching assistant, showcased interactions where students found the LLM beneficial. Participants were presented with two such exemplary interactions, as illustrated in Figure 1B.
@@ 3.2. Approaches to Leverage LLMs in Problem Solving
The manner in which students engage with LLM-based tools is often influenced by the guidance or instructions they receive. The sequence—whether students consult the LLM first or try to solve problems on their own—has implications for both their learning process and the effectiveness of the LLM as a tool. This builds on research in ”when” is it better to provide hints for learning (Razzaq and Heffernan, 2010).
@@ 3.3. Response from Large Language Models
To explore this design space, we conducted two studies - one in an upper-year computer science classroom teaching students database-related concepts and a second in an online controlled setting with crowdworkers solving mathematics problems. Figures 2 and 4 show a high-level overview of both study designs. They are described in detail in the following sections.
@@ 4. Study-1: A formative field study in a CS classroom
This research study was approved by the local ethics board. The data used in this deployment was obtained from a third-year Introduction to Databases course (DB) at a large research-focused North American university in Winter 2023 for Computer Science (CS) students. This course spanned 12 weeks, employing a flipped classroom setup (students were required to complete preparation work: video modules, technical questions, and self-reflections, prior to attending lecture in-person using a learning management system). DB had one course instructor, eleven teaching assistants, and was an upper-year elective course for computer science students. The majority of students were computer science majors, however, some are computer science minors. Although natural classroom limitations impose constraints on achievable statistical power and, as a result, the complexity of questions we can reliably answer with quantitative methods, we used this deployment as a formative one, focusing on both acquiring statistical evidence where it is reliable and exploratory analysis without statistical conclusions to inform larger-scale deployments (both in the virtual lab (Study 2, see Section 5) and in the field (future studies)).
We designed a randomized factorial experiment, for the guidance given to the student, with 2 (G1-List of Suggestions: present vs. absent) x 2 (G2-Example-based instruction: present vs. absent) x 2 (G3-Metacognitive questioning: present vs. absent) x 2 (G4-Solve then refine: present vs. absent) between subjects. The presence of G4 encouraged a Self-First Approach (SFA), while the absence provided more flexibility (LLM-First Approach – LFA) (refer to Section 3). At the beginning of the assignment, students were primed to use LLM-based support tool with the following:
@@ 4.1. LLM-based chat support tool
Each student was given access to a chatbot tutor, based on GPT-3 (described in Section A.1.1), to help them solve assignment problems. The prompt to the model was designed after careful iterations between the researchers of this study, the instructor, and the teaching assistant. The prompt given to the model was the following:
”The following is a conversation with a database instructor. The instructor helps the human solve assignment problems related to database. The instructor never explicitly gives the solution. The instructor also never writes the SQL query. The instructor would only provide brainstorms to possible solutions without providing any SQL statements. This rule should be enforced in the entirety of the conversation. Following are the DDL files to create table given as part of the assignment.”
@@ 4.2. Participants
There were 218 students enrolled in DB, of which a total of 145 students (67%) completed the assignment. When students were asked about their familiarity with Large Language Models (LLMs), 51.03% reported occasional use, 16.55% were regular users, 15.86% had tried them once, and 16.55% had never used them before. On a scale of 1 to 7, the students’ average initial self-efficacy for the topic stood at 4.10 (SD = 1.38).
@@
Verbatim First Assignment Question
The student copies the exact question from the assignment verbatim as the first message to the LLM tutor.
”For each of the following locking protocols (a to d below), assuming that every transaction follows that specific locking protocol…”
28%
First Assignment Question Reworked
The student rephrases or alters the wording of the first question from the assignment, maintaining its original intent.
”Given the following, always obtain an exclusive lock before writing; hold exclusive locks until the end of transaction. No shared locks are ever obtained. Which properties are ensured?”
14%
Unrelated Initial Inquiry/Chatbot Testing
The student’s first query is unrelated to the assignment, often to test the chatbot’s capabilities or driven by general inquisitiveness.
”Hi buddy, recite a Shakespearean sonnet…”
4%
Initial Conceptual Clarification Request
The student seeks to understand specific concepts or parts of the assignment more clearly with their initial query.
”How do I read the Venn diagram for schedules?”
54%
@@ 4.3. Results
We examined the impact of each guidance type (G1–G4) on the nature of the first queries made by students to the LLM tutor. For each guidance type, students were grouped into two categories: those who received the guidance and those who did not. We then analyzed the first queries made by students in each group to see how frequently they fell into each of the established query categories (Figure 3). This allowed us to compare the tendencies of students’ initial queries based on whether they had been exposed to a particular type of guidance. We find that when G1-List of Suggestions, G3-Metacognitive Questioning and G4-Solve then refine with LLM are absent, students tend to ask unrelated questions more, compared to when it is present. On the other hand, the presence of G2-Example based instruction resulted in more students asking unrelated questions, or trying to break the chatbot. G1, G2, and G4 show positive affect on reducing the likelihood of the student copy pasting the exact assignment question, however, G3 showed the opposite effect. The introduction of G3-Metacognitive Questioning resulted in more fraction of students rephrasing the first assignment question, while the absence of G4 had the same effect. For G4-Solve then refine with LLM, we find students who received this particular guidance asked clarification questions more as their first query, in line with what was instructed to them as part of this strategy.
By aligning the types of guidance with the categorized first queries, we were able to identify trends in student interactions. This analysis not only contributed to a deeper understanding of the immediate impact of guidance on student queries but also offered insights into the broader implications of guidance strategies in shaping student learning experiences with LLMs.
@@ 5.1. Participants
Initial self-confidence in solving math problems
1 (1.40%), 2 (3.93%), 3 (8.15%), 4 (18.26%), 5 (28.93%), 6 (22.75%), 7 (16.57%)
@@
Prior use of LLMs
Never (23.88%), Once (20.51%), Occasionally (34.55%), Regularly (21.07%)
Age
18-25 (43.26%), 25-34 (41.57%), 35-44 (10.96%), 45-54 (2.81%), 55-64 (1.12%), 65+ (0.28%)
Gender
Male (47.75%), Female (49.44%), Non-binary / third gender (1.97%), Prefer not to say (0.84%)
Education
Some high school or less (2.25%), High School graduate (10.96%), Some college credit with no degree (23.31%), Associate degree (4.49%), Bachelor’s degree (41.85%), Graduate or Professional degree (14.89%), Prefer not to say (2.25%)
@@ 5.3.1. GPT-3 with Pre-Prompt
”You are a professional K12 math teacher helping students answer math questions.
@@ 6.1. RQ1: Impact of Guidance on Initiating Dialogue with LLM Tutor
6.2. RQ2: Immediate Performance and Long-term Learning with LLMs
Our formative study highlighted a potential design tension in LLM research: the balance between providing immediate assistance and ensuring long-term skill development for students. This fits into the long line of work studying the effects of automated hints on learner performance and learning (Marwan et al., 2019; Corbett and Anderson, 2001; Fossati et al., 2015; Roy Choudhury et al., 2016; Rivers, 2017). An interesting observation in the formative study was the contrast between G4: Solve-Then-Refine (focusing on getting them to solve a problem first, then go to the LLM with questions), vs G1, G2, G3 (which focused on using the LLMs while solving problems). There is suggestive evidence that the presence of G4-Solve-Then-Refine (encouraging self-first approach) may not have helped homework performance as much compared to the other forms of guidance (refer to Table 3). However, the performance in the final exam without any LLM support (considered as a measure of learning) was similar between the different forms of guidance and learner approaches. A recent large-scale controlled study by Kumar et al. (Kumar et al., 2023) showed that when learners received explanations from GPT-4, the long-term benefits were greater for those who attempted problems on their own first before consulting LLM explanations.
@@ 6.6. Future Work & Limitations
There are limitations to the current work. In our formative study, the constraints of small to medium-sized classroom field studies inherently limit the scope of broad statistical generalizations. Additionally, observations from our specific classroom setting of an introduction to database course may not be generalized to other classroom settings and audiences. For example, students with non-computer science backgrounds might interact with LLMs differently and differ in LLM-related trust and confidence. More studies are needed to see whether results from the online controlled setting generalize to learning tasks in other contexts and populations of online learners, who may differ from crowdworkers in their goals and motivation. A larger sample confirmatory study might help to reliably evaluate multiple interactions between student characteristics, experimental variables, achievement, and perception outcomes, which were described as exploratory in the current work. Additionally, our findings could benefit further from a rigorous qualitative analysis of conversations between learners and LLM chatbots, which can help explain the mechanisms behind some of the effects observed in our studies.

### v3

@@ 2.3. Understanding Factors that Influence Learners’ Help-Seeking Decisions
To characterize how learner-LLM interactions may be shaped by important factors that influence learner help-seeking decisions more broadly, we can turn to the educational psychology literature on self-efficacy and perceived task difficulty. The choice of when to consult external tools during problem-solving isn’t exclusive to the realm of LLMs. In broader educational research, students’ decisions to access resources—whether they are textbooks, online forums, peers, or tutors—often come down to an interplay between self-efficacy, perceived task difficulty, and the accessibility of the tool.
@@
The aforementioned considerations have nuanced significance in the context of LLMs. The availability and perceived efficacy of LLMs could encourage more immediate reliance, potentially overshadowing traditional problem-solving approaches. Our paper’s exploration of different student approaches offers insight into these nuanced behaviors in the specific context of LLM interactions.
We factored such considerations into the design of our experiments. We included questions that help us quantify such factors using pre- and post-measures regarding perception of LLMs (confidence in LLMs’ responses, helpfulness, willingness to interact again, and error tolerance) and self-perception (student confidence in their answers, self-confidence for the given topic of the assignment). To measure long-term learning, we assessed their performance on the final exam on the same topic, comprising isomorphic questions to those in the assignment.
@@ 3. Design Considerations for Learner Interaction with LLMs
Understanding how students interact with LLM-based tools becomes imperative as language models continue to find their way into educational settings. In our exploration of learner interactions with LLMs, we identified three critical factors influencing these engagements:
1. Guidance Received
The directions or suggestions that students receive can shape their entire interaction experience. The form and content of this guidance can determine how students approach and utilize LLMs to their advantage (Section 3.1).
Motivation: Recent research highlights the challenges in prompting LLMs for specific tasks (Zamfirescu-Pereira et al., 2023), emphasizing the need for structured guidance to support effective use (Abdellatif et al., 2020; Fiannaca et al., 2023). This is supported by extensive literature that indicates that appropriate guidance can help learners form the correct mental models of tools (Kieras and Bovair, 1984), thus facilitating their proper use. Despite the critical importance of guiding learners in using LLMs, this remains a largely unexplored problem (see Section 2.2). We address this gap by exploring guidance strategies for LLM use.
2. Learner Approach
Learners’ initial inclination to consult the LLM immediately or first wrestle with the problem alone can have implications for their learning process. The strategies they adopt can be influenced by the guidance they have received and can vary significantly in efficacy (Section 3.2).
Motivation: Existing research indicates that individual factors such as self-efficacy, prior exposure, and perceived utility can influence users’ approach when using tools (Moores and Chang, 2009; Vancouver et al., 2002; Glassman et al., 2015). Investigating these approaches in the context of LLM tutors is important, as forming the appropriate reliance and dependency on these tools is necessary for effective learning (Karabenick and Dembo, 2011) (see Section 2.3 for details). Understanding the approaches students take when using LLMs and how guidance can steer these approaches is vital. Some approaches may not be conducive to learning, such as using LLMs to cheat versus using them to seek feedback (Kumar et al., 2023b). Hence, we consider the learner’s approach a crucial design consideration in our study.
3. LLMs’ Response
While guidance and approach set the stage, the actual dialogue with the LLM remains paramount. The quality, clarity, and relevance of the LLMs’ responses can make or break the learning experience (Section 3.3).
Motivation: Research suggests that the accuracy and framing of responses by conversational agents can significantly influence user perception and interaction with the agent (Diederich et al., 2022; Xiao et al., 2007). This is particularly crucial in educational settings where how answers are framed can impact long-term learning outcomes (Weber et al., 2021), even if the immediate effect seems beneficial (e.g., students using LLMs for cheating) (Kumar et al., 2023b). By exploring the steering of LLM responses, we can gain insights into the differences between publicly available LLMs, such as ChatGPT, and those configured by instructors. Additionally, the nature of LLM responses can influence learners’ trust in the LLM and their sense of self-efficacy (Park et al., 2023; Riefle et al., 2022), as direct answers might lead to a realization of lacking knowledge (Yildiz Durak, 2023). These factors prompted us to consider the types of LLM responses as a design element in this work.
@@ 3.1. Guidance Strategies for LLM Engagement
We devised a set of suggestions for the use of LLM chatbots, as depicted in Figure 1A. Given the limited existing research on LLM guidance for students, our suggestions were based on prominent online sources (Pangu, 2023; Miquelino, [n. d.]; Mollick, 2023b). Recognizing the growing classroom trend of prescribed LLM instructions (eth, 2023; Mollick, 2023a), G1 serves as a representative model to evaluate its efficacy against other guidance strategies.
@@
Leveraging examples is an intuitive approach to tool introduction (Lee, 2004). However, the choice of examples plays a crucial role. For our classroom deployment, we took examples of interactions from previous interview studies with students (Kumar et al., 2023a; Xiao et al., [n. d.]). These examples, validated by an instructor and a teaching assistant, showcased interactions where students found the LLM beneficial. Participants were presented with two such exemplary interactions, as illustrated in Figure 1B.
@@ 3.2. Approaches to Leverage LLMs in Problem Solving
The manner in which students interact with LLM-based tools is often influenced by the guidance or instructions they receive. The sequence—whether students consult the LLM first or try to solve problems on their own—has implications for both their learning process and the effectiveness of the LLM as a tool. This builds on research in ”when” is it better to provide hints for learning (Razzaq and Heffernan, 2010).
@@ 3.3. Response from Large Language Models
To explore this design space, we conducted two studies, one in an upper-year computer science classroom teaching students database-related concepts and a second in an online controlled setting with crowdworkers solving mathematics problems. Figures 2 and 4 show a high-level overview of both study designs. They are described in detail in the following sections.
@@ 4. Study-1: A formative field study in a CS classroom
This research study was approved by the local ethics board. The data used in this deployment was obtained from a third-year Introduction to Databases course (DB) at a large research-focused North American university in Winter 2023 for Computer Science (CS) students. This course spanned 12 weeks, employing a flipped classroom setup (students were required to complete preparation work: video modules, technical questions, and self-reflections, prior to attending lectures in person using a learning management system). DB had one course instructor, eleven teaching assistants, and was an upper-year elective course for Computer Science students. The majority of students were Computer Science majors, however, some are Computer Science minors. Although natural classroom limitations impose constraints on achievable statistical power and, as a result, the complexity of questions we can reliably answer with quantitative methods, we used this deployment as a formative one, focusing on both acquiring statistical evidence where it is reliable and exploratory analysis without statistical conclusions to inform larger-scale deployments (both in the virtual lab (Study 2, see Section 5) and in the field (future studies)).
We designed a randomized factorial experiment, for the guidance given to the student, with 2 (G1-List of Suggestions: present vs. absent) x 2 (G2-Example-based instruction: present vs. absent) x 2 (G3-Metacognitive questioning: present vs. absent) x 2 (G4-Solve then refine: present vs. absent) between subjects. The presence of G4 encouraged a Self-First Approach (SFA), while the absence provided more flexibility (LLM-First Approach – LFA) (see Section 3). At the beginning of the assignment, students were primed to use LLM-based support tool with the following:
@@ 4.1. LLM-based chat support tool
Each student was given access to a chatbot tutor, based on GPT-3 (described in Section A.1.1), to help them solve assignment problems. The model prompt was designed after careful iterations between the researchers, the instructor, and the teaching assistant. The prompt given to the model was the following:
{mdframed}
[backgroundcolor=gray!20] ”The following is a conversation with a database instructor. The instructor helps the human solve assignment problems related to database. The instructor never explicitly gives the solution. The instructor also never writes the SQL query. The instructor would only provide brainstorms to possible solutions without providing any SQL statements. This rule should be enforced in the entirety of the conversation. Following are the DDL files to create table given as part of the assignment.”
@@ 4.2. Participants
There were 218 students enrolled in DB, of which a total of 145 students (67%) completed the assignment. When students were asked about their familiarity with Large Language Models (LLMs), 51.03% reported occasional use, 16.55% were regular users, 15.86% had tried them once, and 16.55% had never used them before. On a scale of 1 to 7, the students’ average initial self-efficacy for the topic was 4.10 (SD = 1.38).
@@
Verbatim First Assignment Question The student copies the exact question from the assignment verbatim as the first message to the LLM tutor. ”For each of the following locking protocols (a to d below), assuming that every transaction follows that specific locking protocol…” 28%
First Assignment Question Reworked The student rephrases or alters the wording of the first question from the assignment, maintaining its original intent. ”Given the following, always obtain an exclusive lock before writing; hold exclusive locks until the end of transaction. No shared locks are ever obtained. Which properties are ensured?” 14%
Unrelated Initial Inquiry/Chatbot Testing The student’s first query is unrelated to the assignment, often to test the chatbot’s capabilities or driven by general inquisitiveness. ”Hi buddy, recite a Shakespearean sonnet…” 4%
Initial Conceptual Clarification Request The student seeks to understand specific concepts or parts of the assignment more clearly with their initial query. ”How do I read the Venn diagram for schedules?” 54%
@@ 4.3. Results
We examined the impact of each guidance type (G1–G4) on the nature of the first queries made by students to the LLM tutor. For each guidance type, students were grouped into two categories: those who received the guidance and those who did not. We then analyzed the first queries made by the students in each group to see how often they fell into each of the established query categories (Figure 3). This allowed us to compare the tendencies of the students’ initial queries on the basis of whether they had been exposed to a particular type of guidance. We find that when G1-List of Suggestions, G3-Metacognitive Questioning and G4-Solve then refine with LLM are absent, students tend to ask unrelated questions more, compared to when it is present. On the other hand, the presence of G2-Example based instruction resulted in more students asking unrelated questions, or trying to break the chatbot. G1, G2, and G4 show a positive effect in reducing the likelihood of the student copy-pasting the exact assignment question. However, G3 showed the opposite effect. The introduction of G3-Metacognitive Questioning resulted in a higher percentage of students rephrasing the first assignment question, while the absence of G4 had the same effect. For G4-Solve then refine with LLM, we find that students who received this particular guidance asked clarification questions more as their first query, in line with what was instructed to them as part of this strategy.
By aligning the types of guidance with the classified first queries, we were able to identify trends in student interactions. This analysis not only contributed to a deeper understanding of the immediate impact of guidance on student queries, but also offered insights into the broader implications of guidance strategies in shaping student learning experiences with LLMs.
@@ 5.1. Participants
Initial self-confidence in solving math problems 1 (1.40%), 2 (3.93%), 3 (8.15%), 4 (18.26%), 5 (28.93%), 6 (22.75%), 7 (16.57%)
@@
Prior use of LLMs Never (23.88%), Once (20.51%), Occasionally (34.55%), Regularly (21.07%)
Age 18-25 (43.26%), 25-34 (41.57%), 35-44 (10.96%), 45-54 (2.81%), 55-64 (1.12%), 65+ (0.28%)
Gender Male (47.75%), Female (49.44%), Non-binary / third gender (1.97%), Prefer not to say (0.84%)
Education Some high school or less (2.25%), High School graduate (10.96%), Some college credit with no degree (23.31%), Associate degree (4.49%), Bachelor’s degree (41.85%), Graduate or Professional degree (14.89%), Prefer not to say (2.25%)
@@ 5.3.1. GPT-3 with Pre-Prompt
{mdframed}
[backgroundcolor=gray!20] ”You are a professional K12 math teacher helping students answer math questions.
@@ 6.1. RQ1: Impact of Guidance on Initiating Dialogue with LLM Tutor
6.2. RQ2: Immediate Performance and Potential Long-term Learning with LLMs
Our formative study highlighted a potential design tension in LLM research: the balance between providing immediate assistance and ensuring long-term skill development for students. This fits into the long line of work studying the effects of automated hints on learner performance and learning (Marwan et al., 2019; Corbett and Anderson, 2001; Fossati et al., 2015; Roy Choudhury et al., 2016; Rivers, 2017). An interesting observation in the formative study was the contrast between G4: Solve-Then-Refine (focusing on getting them to solve a problem first, then go to the LLM with questions), vs G1, G2, G3 (which focused on using the LLMs while solving problems). There is suggestive evidence that the presence of G4-Solve-Then-Refine (encouraging self-first approach) may not have helped homework performance as much compared to the other forms of guidance (refer to Table 3). However, the performance in the final exam without any LLM support (considered as a measure of learning) was similar between the different forms of guidance and learner approaches. A recent large-scale controlled study by Kumar et al. (Kumar et al., 2023b) showed that when learners received explanations from GPT-4, the long-term benefits were greater for those who attempted problems on their own first before consulting LLM explanations.
@@ 6.4. RQ5: Learners’ Self-Assessment During the Problem-Solving Task
6.6. Methodological Design Choices for Educational LLM Interventions
Each study helped to address specific aspects of the research questions. With Study-1, our objective was to explore how undergraduate students will interact with LLMs in a natural classroom setting, striving to achieve their course goals. However, this field setting constrained our design space, excluding more risky exploration strategies, potentially disadvantaging some students, and limiting the study sample size. We focused on rich qualitative feedback and interaction data, allowing a detailed examination of how interaction strategies shape student engagement with LLMs and treating quantitative assessments as formative. Consequently, Study-2, conducted with crowdworkers, aimed to build on the results of Study-1 and take a more direct approach to compare the guidance conditions. This allowed us to evaluate individual interaction strategies and the effect of an unprompted LLM, which was infeasible in Study-1 due to ethical reasons.
Although the classroom setting of Study-1 provided useful information on the real-world use of LLMs for learning, the inherent limitation of small sample sizes was a significant challenge (Emerson and Hazlett, 2012). To address this, we used a 2$\times$2$\times$2$\times$2 factorial experiment design, which assessed the presence or absence of four distinct guidance strategies, thus improving the statistical power (Dziak et al., 2012). This contrasted with the approach of Study-2, where the larger pool of participants allowed a broader comparison of strategies and the integration of subjective measures at various intervals (Kittur et al., 2008). These repeated measures could have led to high rates of attrition in the classroom setting. However, the validity of the results of Study-2 may not align with those of Study-1 (Kumar et al., 2023b), where students interacted with LLMs as part of their coursework in a more authentic context.
@@ 6.7. Future Work & Limitations
There are limitations to the current work. Study 1 and 2 differed in various aspects, including study design, participant pool, and subject matter. Although the findings of both studies can be considered complementary, interpreting them together should be done with caution due to the distinct study settings. In our formative study, the constraints of small to medium-sized classroom field studies inherently limit the scope of broad statistical generalizations. Additionally, observations from our specific classroom setting of an introduction to database course may not be generalized to other classroom settings and audiences. For example, students with non-computer science backgrounds might interact with LLMs differently and differ in LLM-related trust and confidence. More studies are needed to see whether results from the online controlled setting generalize to learning tasks in other contexts and populations of online learners, who may differ from crowdworkers in their goals and motivation. A larger sample confirmatory study might help to reliably evaluate multiple interactions between student characteristics, experimental variables, achievement, and perception outcomes, which were described as exploratory in the current work. Additionally, our findings could benefit further from a rigorous qualitative analysis of conversations between learners and LLM chatbots, which can help explain the mechanisms behind some of the effects observed in our studies. Future educational interventions that use LLMs can adopt more adaptive experimental designs such as the Multiphase Optimization Strategy (MOST) or Sequential Multiple Assignment Randomized Trials (SMART) (Collins et al., 2007). These designs offer the flexibility to make dynamic adjustments based on evolving data, which is crucial in educational settings where interaction effects and student responses can vary widely. Using such adaptive strategies and careful design decisions, researchers can better navigate the trade-offs between validity and statistical power, leading to more effective and contextually appropriate LLM interventions for learning.
Deploying guidance strategies for an LLM tutor carries several risks. Students may develop incorrect mental models of the LLM’s capabilities (Kieras and Bovair, 1984), particularly if the guidance (like G1-List of instructions or G2-Examples) does not accurately represent the operational boundaries of the LLM. The choice of examples in G2-Examples is critical; inappropriate examples might mislead students about the applicability or functionality of the LLM, potentially leading to frustration or inappropriate reliance on the LLM tutor (Kulesza et al., 2013; Passi et al., 2024). In addition, introducing complex guidance strategies or excessive information at the beginning might overwhelm the students, leading to cognitive overload (Feldon, 2007). This is particularly a concern with strategies requiring high initial engagement or pre-existing knowledge. The guidance strategies in this study did not specifically account for the diverse needs and backgrounds of all students, such as non-native English speakers (Seale, 2013; Hultgren, 2019). Future work could explore ways to make our proposed guidance strategies personalized for students from varying backgrounds and more accessible to students with special needs.